OpenStudy (anonymous):

Solve 5 over x plus 8 equals 3 over x for x and determine if the solution is extraneous or not. x = 12, non-extraneous x = 12, extraneous x = -12, non-extraneous x = -12, non-extraneous

OpenStudy (anonymous):

@phi

OpenStudy (phi):

can you use the equation editor? It is hard to read.

OpenStudy (anonymous):

sure

OpenStudy (whpalmer4):

\[\frac{5}{x+8}=\frac{3}x\]Cross multiply as a first step: \[5*x = 3*(x+8)\]Expand using the distributive property \[5x = 3x + 24\]Can you do it from there?

OpenStudy (anonymous):

yes its 3

OpenStudy (whpalmer4):

Uh, no. That fact that none of your answers feature the number 3 should be a clue that perhaps you didn't do something right :-)

OpenStudy (anonymous):

but thats not an answer choice and @phi she had put the equation on here

OpenStudy (whpalmer4):

Show me your work...

OpenStudy (anonymous):

i had combine 3x and 5x got 8x=24 divide final answer 3

OpenStudy (whpalmer4):

Explain just how you combined them? \[5x=3x+24\]

OpenStudy (anonymous):

omg my bad its suppose to be 2x=24 which equals to 12 so the answer is A

OpenStudy (whpalmer4):

there you go :-) except you also need to test the solution to see if it is extraneous. If you plug x=12 into the original equation, do you get a number sentence that is true, or false?

OpenStudy (anonymous):

its false so the correct answer is B @whpalmer4

OpenStudy (whpalmer4):

Whoa, hang on, how did you decide it was false? Can you show me your work again?

OpenStudy (whpalmer4):

\[\frac{5}{x+8}\]If we substitute \(x=12\), what does that fraction become?

OpenStudy (anonymous):

i replace x with 12

OpenStudy (whpalmer4):

what are the two fractions you get after you replace \(x\) with \(12\)?

OpenStudy (anonymous):

i was looking at the wrong problem

OpenStudy (anonymous):

do you think you could help me with one more

OpenStudy (whpalmer4):

One of the most valuable things learned in math class is working carefully, with attention to detail... Sure, what's the other question?

OpenStudy (anonymous):

What are the possible number of positive, negative, and complex zeros of f(x) = -3x4 - 5x3 - x2 - 8x + 4? Positive: 2 or 0; Negative: 2 or 0; Complex: 4 or 2 or 0 Positive: 1; Negative: 3 or 1; Complex: 2 or 0 Positive: 3 or 1; Negative: 1; Complex: 2 or 0 Positive: 4 or 2 or 0; Negative: 2 or 0; Complex: 4 or 2 or 0

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

i know there is only one sign change

OpenStudy (whpalmer4):

I get a notification when you respond on a post where I've commented, it isn't necessary to tag me again and clutter up my life with additional notifications, thanks... Okay, you've got the polynomial written in the right order (descending powers of the exponent), and I agree, there is one sign change in \(f(x)\). That means we have 1 positive real root, or 1 negative real root or 1 complex root (you tell me which)?

OpenStudy (anonymous):

oh well my bad i didnt know because i dont get a notification when you respond back

OpenStudy (anonymous):

1 positive real root

OpenStudy (whpalmer4):

I think the system assumes that you'll be watching your own problem intently :-) Anyhow, if I haven't responded within a few minutes, then feel free to tag me, just don't do it immediately. Yes, 1 positive real root is correct. What's the next step in the procedure?

OpenStudy (anonymous):

ok i got you:) and next you should try to find the negative real root right

OpenStudy (whpalmer4):

Try to find the count of the possible negative real root(s), that is. What do you do to do that?

OpenStudy (anonymous):

look at a number that is negative right

OpenStudy (whpalmer4):

No, you need to count the possible negative roots. You change \(f(x)\) to \(f(-x)\) and do the sign change count again. That gives you your possible negative real roots.

OpenStudy (anonymous):

so would it be 3

OpenStudy (whpalmer4):

\[f(x) =-3x^4 - 5x^3 - x^2 - 8x + 4\] \[f(-x) = -3x^4+5x^3-x^2+8x+4\] 3 sign changes in \(f(-x)\) so we have either 3 negative real roots or 1 negative real root and 2 complex roots. How many roots will we have in total?

OpenStudy (anonymous):

4

OpenStudy (whpalmer4):

Right. So what are the possible scenarios for the composition of the roots of this equation, if we don't know what they actually are?

OpenStudy (anonymous):

B right

OpenStudy (whpalmer4):

No, I want you to spell out the combinations in your own words...

OpenStudy (anonymous):

ok so there is 3 or 1 negative roots and one positive root

OpenStudy (whpalmer4):

Hey, OS, don't mess with what I type :-( The whole story: We can have: 1 positive real root, 3 negative real roots, and 0 complex roots, OR 1 positive real root, 1 negative real root, and 2 complex roots in a conjugate pair conjugate pair means they can be written as \(a\pm bi\) where \(i = \sqrt{-1}\)

OpenStudy (whpalmer4):

we know the complex roots are in a conjugate pair because we only have real numbers as the coefficients of this polynomial. To get the complex roots to not contribute any complex coefficients, they have to get combined as the difference of squares: \[(a+bi)(a-bi) = a^2 -abi + abi -b^2i^2 = a^2-b^2i^2=a^2-b^2(-1) = a^2+b^2\]Look ma, no \(i\)!